### v1

# Vulnerabilities that arise from poor governance in Distributed Ledger Technologies

Abstract.
Current implementations of governance in Distributed Ledger Technologies leave them susceptible to a number of attacks. We survey the state of the art of Distributed Ledger Technologies’ (DLTs) governance protocols and work carried out to systematise good governance properties in the context of DLTs. We then select the most appropriate taxonomy of good governance properties and point to formal security notions that good governance protocols should satisfy. We point practitioners to existing solutions to deliver them, where possible. Furthermore, we outline a number of vulnerabilities that arise in the absence of good governance properties. We call on the research community and DLT research practitioners to prioritise delivering these good governance properties and continue to develop tools to do so, to avoid attacks to DLT protocols that exploit their poor governance models.
Governance, Distributed Ledger Technologies, Privacy, Voting
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Human computer interaction (HCI)$c$s: Security and privacy
1. Introduction
Distributed Ledger Technologies (DLTs) have enabled a paradigm shift in how users interact and transact online. DLTs are comprised of users, that interact with the DLT network and receive some service from it, and service providers (miners and node operators) that maintain the DLT network by running its protocol and providing service to its users. DLTs do not exist in a single server, but rather run on a web of connected nodes, operated by the service providers, which communicate following a protocol. A protocol is simply a set of instructions that nodes execute to enable the functioning of the DLT. DLTs are run through code, and maintainable code requires updates. DLTs must have a mechanism through which protocols are updated, bugs are patched and the service is improved. This component is crucial to ensure DLTs adapt, develop and improve in a sustainable manner, as well as providing a much needed protection against threats. Without this ability to mutate, vulnerabilities cannot be patched. Since all nodes must agree to use the same protocol, a DLT must rely on some mechanism utilising the underlying consensus algorithm to decide which protocol updates will take place and when. Governance is an all-encompassing umbrella term that captures this ability: the decision-making process to change a DLT protocol.
Governance is (one of) the Achilles’ heel of DLTs. It is central to the promise of decentralisation and it is difficult to untangle. There are many different research lenses from which it can be approached, and often they cannot be divorced from one another. Governance is at the intersection of numerous research areas: legal, software, political, institutional, philosophical, cryptographic, social choice, economic and financial are just some examples. Extensive literature exists outlining parallels with DLT governance and other organisational models such as DAOs, or outlining desirable properties, however, we find it particularly challenging to extract tangible solutions from existing work. Indeed, there is literature outlining what are desireable properties of governance, but little guidance on how to actually achieve and implement said criteria and what are the practical risks when these are not satisfied. We note that there has been a recent upsurge in the research of DAO governance, and we argue that parallels may be drawn from DAO and DLT governance. This follows naturally from the fact that DAOs are often built on DLTs.
Our research focus is on security and its implications in different aspects of the governance protocol. We point practitioners on what existing solutions are available to achieve certain governance properties, and what are the security risks of failing to do so. We outline which desirable properties are currently mutually exclusive, and which need not be.
2. Related Work
A comprehensive systematization of knowledge on good governance properties can be found in (Kiayias and Lazos, 2022) and (Liu et al., 2023). The authors in (Kiayias and Lazos, 2022) focus on governance properties, whereas (Liu et al., 2023) answer questions on what governance is, who is involved, when is it used and why. The work in (Kiayias and Lazos, 2022) serves as a starting place for the work we present. Namely, we find their taxonomy of governance properties the most clear, tangible and concise thus far. However, they explicitly note that they do not provide specific techniques, algorithms or mechanisms to achieve good governance properties.
The authors in (Khan et al., 2020) consider blockchain governance in the context of IT governance, and study different decisional outcomes using the Nash equilibrium. Other research efforts such as (Pelt et al., 2021), (Beck et al., 2018) and (De Filippi and McMullen, 2018) aim to define governance and its components, although none delve deeply into tangible definitions of good governance properties.
In terms of vulnerabilities that may arise as a consequence of poor governance in DLTs, (Schneider, 2022) claims that economic incentives alone do not suffice to enable good governance, and makes a case that well studied social choice and politics findings should be applied to governance in DLTs. The work in (Benedict, 2019) outlines a number of challenges regarding DLT governance. Namely, it notes how poor governance models can cause oligarchies in DLTs, how incentives are crucial in ensuring accountability and notes how decision rights should be clearly defined in the governance protocol. The authors in (De Filippi and Loveluck, 2016) identify that governance in Bitcoin is not as decentralised and code-driven as expected, and uncovers that in fact the decision-making power is highly technocratic and determined by a small minority of code developers. None of the aforementioned works address the formal security and privacy notions necessary in the governance protocol, or the vulnerabilities that arise in the absence of these. We find a substantial overlap in the study of governance vulnerabilities in DAOs and DLTs. The work in (Austgen et al., 2023) and (Daian et al., 2018) outline how vote buying cartels may be automated as a consequence of poor governance models in DAOs. A noteworthy body of research is carried out in (Tan et al., 2023), where the authors note the importance of classical security notions for e-voting should also hold in DAO governance mechanisms. We agree. However, they do not provide guidance on how to achieve these primitives. In summary, some recent works have emerged outlining vulnerabilities that arise as a consequence of poor governance models in DAOs, but less so regarding DLT governance. Some cautionary warnings have been raised regarding DLT governance, but we find no work that explores the suite of possible attacks that may arise in DLTs in the absence of standard good governance properties.
3. Governance Components
In this work, we refer to governance as the decision-making processes to enable the functioning and updating of a DLT protocol. Before proceeding with necessary criteria that good governance should satisfy, we first outline what are the crucial components of any governance protocol:
  • The voters.
  • The votes.
  • The vote aggregation algorithm.
  • The election officials. Note that these are participants in charge of tallying the votes, maintaining the voting platform and publishing results.
  • The bulletin board. This is the digital record of cast ballots.
  • Stakeholders: they may or may not be allowed to vote, but they are in some way affected by the outcome of an election. This includes participants in the DLT that are in charge of implementing the election outcome.
  • The voting mechanism. This describes how votes are tallied to return a winner.
  • The platform to vote.
DLTs allow for novel ways to implement and deliver some of these components. This opens a wide range of opportunities, but like any new and unexplored field, may come with landmines. To understand these risks, we begin with reiterating the taxonomy of governance properties first introduced in (Kiayias and Lazos, 2022). The reader should note the subtle distinction: the following are properties of the aforementioned governance components.
4. Governance Properties
4.0.1. Participation
First we begin by considering who may participate in the decision making process of DLTs. These actors can be classified into the following set of agents: users, service providers, proposers and decision-makers.
Definition 4.1 (User).
A user is a participant of the DLT network that interacts with the DLT to receive a service from it.
Definition 4.2 (Service providers).
A service provider runs a set of agreed upon instructions (protocol) in their own machine that enables the functioning of the DLT.
Definition 4.3 (Decision-makers).
A decision-maker is an agent that has the power to vote on a new proposal.
Definition 4.4 (Proposer).
A proposer is an agent that can put forth a protocol update which will be subject to vote.
Each of these sets of agents may overlap, service providers may also be proposers, users may also be decision-makers and so forth. The intersection of these sets is application specific to each DLT. There are two questions that must be addressed with regards to participation: suffrage and distribution of power. The former addresses who is allowed to vote, and the latter, how power is distributed amongst the voters.
We advocate for equal suffrage. This property removes the possibility of weighted votes, thereby enforcing equal voting power distribution.
This property may be too strict in some contexts. Some DLTs may wish to grant voting rights proportional to reputation or trustworthiness. However, we strongly discourage practitioners from distributing voting power proportionally to wealth.
In summary, two important properties of governance participation are:
  1. P1
    Equal suffrage.
  2. P2
    Non-purchasable voting power distribution.
Note that P2 is a weaker notion of P1. P1 states each voter should receive one vote and all votes should be of equal weight. P2 allows for votes of different weights, as long as the weight cannot be purchased.
Possible implementations: Property P1 advocates for a one-person-one-vote mechanism to distribute voting power. This can be easily achieved in permissioned DLTs, since each participant’s identity is known. In permissionless DLTs, this is more challenging to achieve if a decentralised structure is to be retained [^1] It can be argued that some permissionless DLTs are already not truly decentralised if the majority of voting power is held by two or three entities and participation rights are assigned proportional to wealth. See Figure 1 for a distribution of hashrate in Bitcoin (regulated through Proof-of-Work) and Figure 2 for the staking pool distribution of Ethereum (regulated through Proof-of-Stake)..
Figure 1. Bitcoin’s hashrate distribution for the year of 2024 to 2024. Source: Blockchain.com
Figure 2. Ethereum’s staking pool distribution for the first 6 months of 2024. Source: Dune.com
This is because it is non-trivial to distribute unique identity credentials in a decentralised manner, and in a way in which a single user cannot easily obtain multiple identity credentials. Indeed, an ongoing research effort is dedicated to answering this question (Dunphy et al., 2018), (Goodell and Aste, 2019). The Proof of Humanity project aims to address this challenge. It is a peer-to-peer protocol that issues unique identity credentials to its participants(PoH, 2021). They each create an identity, and it is then approved by other validated participants in the protocol. It offers the benefit that it is a more decentralised solution (it distributes trust), but participants must wait for others to vouch for their profile, and it is built on Ethereum so it is only available on their ecosystem. Bridges are protocols that allow for the transfer of data or transactions across different blockchains, and may be leveraged in this case to enable PoH on other compatible blockchains (Lee et al., 2023).
Technologies such as Semaphore, which is an Ethereum based zero-knowledge protocol, allows members to cast a vote without revealing their identity, whilst demonstrating they belong to a group of valid voters (Semaphore projects, 2022). In this manner, voters can only participate if they prove in zero-knowledge that they own a valid, unique voter credential, and can vote without having to reveal it. It is possible to assign one person one vote in permissionless DLTs by combining a protocol such as Proof of Humanity with Semaphore. Proof of Humanity ensures only approved (real) individuals receive a unique identity credential. Then, voters can prove they own such a credential in zero-knowledge, meaning they do not need to publicly disclose their credential, through Semaphore. The work in (Gangemi and Kharman, 2023) proposes combining these two protocols in the context of resolving disputes is Web 3 platforms. We note both Proof of Humanity and Semaphore are only compatible with Ethereum. We encourage DLT engineers and researchers to prioritise delivering these functionalities across varying DLT architectures.
In summary, one-person-one-vote is a fair and democratic way of allocating voting rights but it will come at the expense of re-introducing some form of centralisation in permissionless DLTs to regulate who is considered a valid voter, and ensuring they only receive one unique identity credential.
4.0.2. 4.2 - Incentives
Many DLT protocols assume their governance system is secure due to the financial incentives offered to their participants. Unfortunately, in a number of cases these governance protocols are not clearly defined, documentation is sparse, outdated or spread across multiple sources (Foundation, [n. d.]), (IOTA, 2023), (IOTA governance forum, [n. d.]). Some protocols are so complex it is non-trivial to outline a clear mathematical model of how all of the incentives interact and what the net best strategy for an agent would be (Polkadot, 2023). As a result, undesirable scenarios arise whereby agents are inadvertently incentivised to undermine the network (Zook and Blankenship, 2018). There are two crucial incentives that must be clearly defined and identifiable in the governance protocol:
  1. I1
    Incentive to vote.
  2. I2
    Incentive to make useful and beneficial proposals.
Property I1 encapsulates whether the voter has an incentive to participate honestly in the voting protocol, and I2 ensures there is no incentive to make undermining proposals.
We then consider two more necessary properties:
  1. I3
    Distributed liability: risk is appropriately distributed amongst the governance participants.
  2. I4
    Binding outcome: the outcome of the election must be implemented, provided the outcome is valid.
I3 is necessary to encourage the implementation of valid election outcomes. If service providers have no liability for not complying, I4 is hard to enforce. Service providers may simply choose to not implement the outcome, without incurring on any cost for doing so. DLT practitioners must select an appropriate function to distribute risk, which is context specific to their own application. We highlight, it does not suffice with distributing liability amongst agents: it must be clear and traceable how this liability is distributed as well. Election outcomes must also be binding, otherwise the governance mechanism is simply a polling or referendum mechanism.
4.0.3. 4.3 - Voting Mechanism
Social choice theory has long studied the different properties that vote aggregation algorithms offer. Arrow famously demonstrated that no ordinal [^2] Ordinal voting schemes only capture preference, whereas cardinal schemes reflect how much more one candidate is preferred over another. voting scheme can simultaneously satisfy three seemingly simple and desirable properties:
  1. V1
    Independence of Irrelevant Alternatives: If A is preferred over B, then introducing an irrelevant option C should not affect the outcome.
  2. V2
    Pareto efficient: if the majority of voters prefer A over B, the voting system should output A over B.
  3. V3
    Non-dictatorial: there is no one voter that can single-handedly determine the election outcome.
(Arrow, 1950). The Gibbard–Satterthwaite theorem further outlines that no ordinal voting scheme is simultaneously non-dictatorial and robust against tactical voting (Satterthwaite, 1975). In other words, that it does not satisfy V3 and V4 where V4 is defined as:
  1. V4
    Strategy-proofness: The best way for a voter to achieve their preferred outcome is to vote according to said preference.
This does not bode well for ordinal voting schemes, but none of these results apply to cardinal voting schemes. A caveat with cardinal voting schemes is that it is important to find a meaningful function to allocate value to preferences. One example is Quadratic Voting (Lalley et al., 2016), whereby each extra vote for a candidate is quadratically more expensive for a voter. Another class of voting mechanisms are those that are probabilistic. An example is the serial random dictator model or Maximum Entropy Voting (Sewell et al., 2009). Probabilistic voting mechanisms offer desirable properties such as V1, V2 and V4 (or slightly weaker notions of the latter two), and in some instances even V3 (Sewell et al., 2009), (Bade, 2015), (Kosheleva et al., 2021), (Bade, 2020). We remind DLT designers of the classical desirable properties of voting schemes, although these can be side-stepped for slightly weaker notions in order to achieve a better compromise. Nonetheless, we strongly advocate for cardinal (provided a meaningful utility function to allocate preference intensity exists) or probabilistic schemes.
4.0.4. 4.4 - Governance Liveness
In the context of blockchain, liveness is the property that ensures nodes in the network are available to exchange information and eventually achieve consensus (Alpern and Schneider, 1985). In the context of governance we define the property of liveness as follows:
  • L1
    liveness in governance: the governance process eventually terminates and returns a valid outcome.
This property is crucial to ensure the system cannot be attacked by refusing to output an outcome. This is especially important in states of exception, for example in the case where a major vulnerability has been discovered and must be patched.
4.0.5. 4.5 - Security
In this section we present a non-exhaustive survey of formal notions of security for online voting schemes, and existing implementations that satisfy these. We highlight that there is no academic consensus as to what the formal, cryptographic definition of these notions should be. DLT practitioners should select the primitives that offer the security level necessary or desired for their application.
Let us proceed by outlining the core notions of secure voting schemes. They must ensure the following properties are satisfied (Jardí-Cedó et al., 2012), (Neumann, 1993), (Petride, 2002):
  • S1
    Ballot Secrecy: it should not be possible to determine how a voter voted from their ballot.
  • S2
    Verifiability: voters should be able to assert that the election outcome reflects the will of the voters.
  • S3
    Coercion resistance: a voter can not comply with the instructions of an adversary on how to vote, and this behaviour cannot be detected by the adversary.
  • S4
    Censorship resistance: it should not be possible to suppress a valid voter’s ballot.
Ballot Secrecy: The quest to formally define S1 is non-trivial. A recent survey of formal definitions of Ballot Secrecy is available in (Smyth, 2021). In it, the author also presents a formal definition of ballot secrecy. It is the state-of-the-art and it accounts for an adversary that has the power to intercept ballots as they are cast. In the same work, Smyth also proves that a variant of Helios voting scheme, namely a variant of Helios 3.1.4 satisfies ballot secrecy. This variant of Helios is known as Helios’16 and can be found in (Smyth et al., 2017).
Possible implementations: A number of voting scheme implementations exist that satisfy notions of ballot secrecy. Well studied proposals include Helios(Adida, 2008), JCJ (Juels et al., 2005), Civtas (Clarkson et al., 2008) and Athena (Smyth, 2019). All of these protocols offer encrypted ballots, but not all provably satisfy the same notions of ballot secrecy, some have weaker security (earlier versions of Helios), others stronger, (JCJ and Civtas) but pay a price for it with an increased computational complexity. We note that none of these schemes have been implemented on-chain, although there are numerous voting smart contracts available. Whilst other on-chain voting schemes exist, most do not provide encrypted ballots. We find VoteCoin (Scholtz, 2022) and MACI (Buterin et al., 2023) are the only on-chain voting schemes that provide encrypted ballots. Unfortunately, VoteCoin provides encrypted ballots only while the election is occurring. This forgoes secrecy once the election is over, enabling coercers to check if their victim complied. The same can be said about the work in (Košt’ál et al., 2019). In MACI, votes remain encrypted, but the coordinator tasked with tallying individually decrypts ballots. If these are intercepted or if the coordinator is not trustworthy, votes may be easily revealed. Note that other protocols exist that provide anonymous voting, whereby the vote is public but the identity of the voter is hidden. Examples include Snapshot (Labs, 2023) and Vocdoni (Vocdoni, 2021). This is not congruent to secret ballots and should not be considered an appropriate security measure especially in DLT implementations, where identities are at best pseudonymous.
In summary, DLT practitioners have a number of options to ensure their governance protocol satisfies notions of ballot secrecy. They may run off-the-shelf protocols such as Helios’16, Civtas, Athena or others previously mentioned. These can be run off-chain, and the results may be notarised on-chain leveraging zero-knowledge proofs. These can be used to demonstrate the election result was correctly computed. Alternatively, DLT practitioners may implement these voting schemes on-chain, although this option will incur greater time and cost. We discourage from following this path. The aforementioned voting schemes have been well studied and audited. Implementing these on-chain would require security audits and verifying that no bugs arise or inaccuracies appeared. Alternatively, if on-chain voting is essential, practitioners may settle for options such as MACI, VoteCoin or Vocdoni which offer encrypted ballots, but these options are insecure and neither meet formal notions of ballot secrecy.
Verifiability: An informal notion of S2 in electronic voting schemes is end-to-end verifiability. If this is satisfied, a voter should be able to verify their ballot was:
  • cast as intended,
  • recorded as cast,
  • tallied as recorded.
(Benaloh et al., 2015). Another informal notion of verifiability is universal and individual verifiability. The authors in (Kremer et al., 2010) state verfiability can be decomposed into:
  • Individual verifiability: voters can check that their own ballots are recorded.
  • Universal verifiability: anyone can check that the tally of recorded ballots is computed properly.
  • Eligibility verifiability: anyone can check that each tallied vote was cast by an authorized voter.
The work in (Smyth et al., 2015) then provides new formal cryptographic definitions for the three aforementioned.
Whilst the informal notion of verifiability is rather intuitive, formalising such notion has been an arduous endeavour. For a comprehensive review of the literature and a proposed formal notion of verifiability, the reader may refer to the work in (Cortier et al., 2016).
Possible implementations: regarding available implementations, all e-voting schemes satisfy notions of verifiability (some weaker than others). It becomes a matter of deciding the level of security and the trust assumptions one can make in their system, to pick the right verifiability formalisation. Achieving verifiability alone is trivial, as can be seen from most blockchain voting implementations: simply publicly revealing votes yields verifiable voting schemes. The challenge lies in achieving both verifiability and secrecy, without sacrificing one in favour of the other. We stress: practitioners do not need to compromise, achieving both is indeed possible, and is in fact necessary. The protocols outlined in 4.0.5 provide verifiability as well as ballot secrecy (granted that each adhere to different formal notions of both ballot secrecy and verifiability).
Coercion Resistance: The work in (Juels et al., 2005) introduced a paradigm shift in the concept of coercion-resistance. Prior to it, Benaloh and Tunistra introduce the notion of receipt-freeness as means to ensure voters cannot prove to a thrid party how they voted (Benaloh and Tuinstra, 1994). However, the authors in (Juels et al., 2005) consider a much more powerful adversary, one that may ‘demand of coerced voters that they vote in a particular manner, abstain from voting, or even disclose their secret keys’. They provide a formal notion of coercion resistance, and an election protocol that provably satisfies said notion. Note that this definition of coercion resistance partly captures censorship resistance, as the adversary may demand of the voter to abstain. However, it does not consider an adversary that may suppress a ballot from the tallying phase. A recent survey of coercion resistant voting schemes can be found in (Battagliola et al., 2024) in section 6.5.1, and their respective time complexities is shown in Table 6.2. None of these schemes are implemented on-chain. The authors in (Spadafora et al., 2020) propose a coercion resistant voting protocol using blockchain to cast votes. They also propose a security notion of Vote Indistinguishability (VI), which is congruent to secret ballots, and show their voting protocol is VI secure.
Possible implementations: for on-chain voting, DLT practitioners may implement (Spadafora et al., 2020). Alternatively, one can select any of the aforementioned e-voting schemes that are coercion resistant (for example (Clarkson et al., 2008)), run them off-chain and notarise the result on-chain leveraging zero-knowledge proofs. Finally, a number of other coercion-resistant schemes have been proposed but not implemented. These may be implemented on-chain if the DLT of choice allows for smart contracts in layer 1, or similarly, implement them off-chain and notarise the result on-chain.
5. Missing properties
This section outlines the number of currently disregarded good governance properties. Many DLT governance protocols are currently failing to meet standards of good governance in their implementations. Some of these include: 1. Binding Elections: Many governance protocols host their voting on off-chain platforms, meaning the results are not binding. Similarly, some schemes allow participants to vote that are not responsible for implementing the results.
2. Censorship resistance: It must not be possible to suppress voter’s votes.
3. Incentive alignment: The governance protocol must not create active incentives to attack it. In other words, if voters act following a strategy that maximises their benefit, this strategy should not be one that actively undermines fairness, security and decentralisation in the governance protocol.
4. Non-purchasable voting power: Voting rights should not be distributed proportionally to wealth.
5. Fee-less voting: Voters should not be required to pay to vote.
6. Verifiability: Voters should be able to verify the election outcome reflects their votes.
7. Ballot Secrecy: Votes must remain secret. Ballot secrecy should not come at the expense of verifiability.
The absence of these governance properties results in a number of possible exploits. We further outline these in the subsequent section. In current DLTs, achieving all the above properties is not possible, since some are mutually exclusive. We encourage practitioners to consider the context specific security and performance needs of their DLT and select a compromise appropriately. We proceed by outlining which properties are mutually exclusive and why that is the case.
Binding elections and fee-less voting: To ensure that election outcomes must be implemented on the DLT, these must be held on-chain. However, on-chain voting sacrifices fee-less votes, because to cast a vote on-chain one must pay transaction fees. On-chain voting in DLTs also currently sacrifices ballot secrecy in favour of verifiability, although this need not be the case.
Binding elections and censorship resistance: On-chain voting cannot guarantee censorship resistance. However, this concern is application specific, as some DLTs may offer increased levels of censorship resistance due to their protocol design and network distribution. This is also dependant on how strong a notion of censorship resistance DLT practitioners wish to provide.
Non-purchasable voting power and permissionless access to the DLT: This is currently the most prevalent issue in DLT governance. DLTs were designed to avoid having to trust central authorities in order to deliver a service. In order to achieve this whilst being protected against resource attacks, they must still offer some protection against Sybil attacks. In other words: because DLTs aim to be decentralised, they must avoid scenarios where any actor can access the network and attack it. Therefore, access to the DLT must be regulated. Currently, to avoid doing this by having to again rely on a centralised trust authority to grant access credentials, access is regulated through a proxy for wealth. The wealthier amass more voting power as a consequence. De-facto, this re-introduces centralisation in the DLT. We advocate for access control mechanisms that do not prioritise the wealthier, but are based on other criteria. An example is the Tree-Proof-of-Position mechanism presented in (Kharman et al., 2023). This work presents a decentralised class of algorithms that allow agents to prove they are in a given position in a collaborative manner. A scheme like this may be used to regulate access to DLTs, such that only agents with a valid proof of position may be granted participation rights.
Secret ballots and verifiability: Current implementations of DLT governance also forgo ballot secrecy in favour of verifiability. Votes are cast publicly on-chain, or are publicly decrypted at the end of an election. A number of e-voting schemes have been developed to achieve both. Examples include: (Adida, 2008), (Ryan et al., 2009), (Clarkson et al., 2008), and even proposals that address coercion resistance such as (Juels et al., 2005).
6. Vulnerabilities
In this section we outline the existing vulnerabilities that arise as a consequence of poor governance models. We highlight the intersection between DAO and DLT governance: both share similarities in their practical implementations. Indeed, this follows from the fact that many DAO projects are built on DLTs.
6.1. Exploits leveraging public votes
The right to a private vote is recognized as a human right as per Article 21.3 of the Universal Declaration of Human Rights (UN General Assembly, 1948) in 1948. In the absence of it, voters may be coerced, bribed or sell their votes. Secret ballots are congruent to freedom, they ensure voters can express their will free of external influences and pressures. This is no different in DLT governance. The absence of secret ballots is responsible for a suite of vulnerabilities that we now proceed to outline.
6.1.1. 1. Voter Coercion
When ballots are public, everyone can know how a voter voted. This provides verifiability: everyone can check that the election outcome is correct and that only those allowed to vote voted. However, this forgoes freedom. Free will cannot be protected when a voter faces the reality of not abiding with their coercer’s threats (Lever, 2007). The coercer can easily check if their victim complied. Voters may chose to sell their votes even in the absence of coercion. In some cases, DLTs provide a unique set of conditions where the most benefitial strategy for some voters is to sell their vote. Other works have identified this issue in DAO governance (Daian et al., 2018), (Austgen et al., 2023). We highlight that this problem is prevalent in DLT governance too.
Currently, most DLTs have informal channels of discussion where participants debate on whether the proposal will be implemented, and from what time on-wards it will be adopted if it has received enough support. These channels range from Slack, GitHub or Discord, to dedicated voting portals such as Snapshot and Tally, and often involve ‘rough consensus’.
We stress, at current time of writing, none of these solutions offer secret ballots. No existing solution can be shown to satisfy formal notions of ballot secrecy, because the ledger is public. Some have attempted to provide increased levels of privacy, but this privacy is short lived, and often compromised to achieve verifiability. An example is Snapshot’s recently added feature of ‘shielded voting’, wherein votes remain private only until the end of the election. Then, ‘when voting ends, all votes are revealed - as well as who voted for what’  [^3] Snapshot adds ‘Shielded Voting’. This still enables voter coercion and vote selling.
6.1.2. 2. Disincentive to vote against whales
Both users and miners vote with their feet on a DLT. When a proposal arises for a protocol update, both must agree for it to be implemented. Users may favour a certain proposal, but miners can refuse to service transactions following it. Similarly, if miners favour a proposal that users disagree with, the users may simply stop sending transactions on that chain, and miners will earn no fees. Let us consider a voter with much larger voting power than the rest a whale. An example could be a user with a large share of tokens or a collective with high mining power (a large mining pool) in the DLT. When a whale votes for a certain proposal, other agents are incentivised to follow suit. Otherwise a fork may arise, and eventually one of the two will outgrow the other and become the main chain. Miners processing transactions on the losing fork will not receive their reward. Consequently users’ transactions will not get approved, and thus they too will eventually leave the fork. Because whales hold greater voting power, the fork they vote for will be the winning one with greater likelihood. Less influential voters will minimise their chance of losses by following the behaviour of whales, which is public, on-chain. This creates a herding effect, and a disincentive to vote true to one’s beliefs, if these oppose those of a whale’s.
For a voter in this situation, their only non-losing strategy is to abstain, but they are guaranteed a reward when they sell their vote. Given that votes are public, this sale can be verifiable: the vote buyer can check that they complied. Indeed, (Austgen et al., 2023) and (Daian et al., 2018) show how vote buying can be automated through a DAO cartel. Their prototype for vote buying code is designed to buy DAO votes, although there is no impediment to using it to buy DLT votes.
6.1.3. 3. Voter Information Asymmetry
When ballots are public, the first voter casting a vote has no information on what the election outcome will be, whereas the last voter to cast their vote knows the election outcome. Voters are unequal: the later you vote the more information you obtain regarding the election result. This allows more powerful voters to withhold their vote until they can guarantee that it is sufficient to sway the election result in their favour.
‘Pump and dumps’ become easy: a proposal may be submitted that stands to inflate token prices. A whale holding many of these tokens may informally communicate that they will vote for this proposal (i.e.: via Twitter, Discord or other similar channels). Less influential voters begin to cast their votes in agreement (we explain above how there is no incentive to vote against the views of a whale). Token prices begin to surge, as it is expected for this proposal to pass. Participants rush to buy more of these tokens. All the meanwhile the whale has not cast their vote yet. Just before the end of the election, the whale sells their tokens, casts their vote, and thus crashes the token price. They have made a profit by cashing out, exploiting the voter information asymmetry that public votes provide, tanking the token price in the process. The public votes allows all participants in the DLT to behave in agreement, creating a herding effect. It is this behaviour that the whale can then exploit for their own profit. Private votes prevent this scenario.
Another simpler exploit consists of an attacker purchasing enough voting power to approve a malicious proposal in their favour. Public voting facilitates this attack: the attacker knows exactly how much voting power they need to acquire to overturn an election in their favour. These attacks have been witnessed in DAOs a number of times  [^4] Build DAO’s hostile governance takeover attack, Feb 2022,  [^5] Beanstalk cryptocurrency project robbed after hacker votes to send themselves \$182 million,  [^6] Sanctioned Tornado Cash DAO governance heisted by hacker, and are plausible in DLTs too. As shown in (Wang et al., 2020) and (Gencer et al., 2018), miners’ voting power (hash rate distribution) is highly centralized. In some cases, the voting power held by one agent can suffice to fork a network. In others, malicious users may leverage flash loans [^7] Loans written in smart contracts that enable participants to quickly borrow funds without the need for collateral (Chainalysis, 2023). to rapidly obtain sufficient tokens, submit a proposal to their benefit, cash out and repay the loan, whilst still making a profit. Miners are incentivised to service this proposal: they stand to gain high transaction fees. Indeed, the behaviour of whales is known to affect the price of a token (Manahov, 2022): they have the power to send large scale transactions that may incentivize miners to act against the consensus protocol (Liao and Katz, 2017), (Daian et al., 2020).
6.2. Exploit leveraging non-binding elections
In a DLT context, we consider a binding election one in which the outcome of the election is enforced into the DLT.
We note that, although informal voting channels may exist, these are in fact nothing other than a polling mechanism, because the only binding vote on a DLT protocol update is in fact the transactions sent by users and approved by miners on-chain, following either one protocol or another. All other off-chain so-called voting platforms are operating under the assumption that miners and users will behave on-chain in the same way that they expressed they would off-chain. The New York Agreement is a clear example of how off-chain voting is not binding and how this assumption may not always hold [^8] Bitcoin’s New York Agreement (Hunter, 2022). Anyone can vote, but if the agents responsible for implementing the election outcome disagree with the outcome, they have no obligation to implement it, even if the election was fair, free and correct. This is a vulnerability, as stakeholders may subvert an election outcome by refusing to comply. To prevent this, DLT governance protocols must be binding. To ensure this, elections must be carried out on-chain. On-chain voting brings a suite of challenges with it, which we outline in section 6.3.2.
6.3. Exploits leveraging purchasable voting power
6.3.1. 1. Plutocracy
1-dollar-1-vote mechanisms and any variant of these where voting power may be purchased will converge towards granting power to the most wealthy. Any governance mechanism that allows users to purchase voting power will inherently converge to plutocracy and centralise over time (Reijers et al., 2021). Whilst this issue is well known, DLTs continue to implement token based voting, for lack of a better alternative for a decentralised Sybil-protection system. DLT practitioners must compromise and choose which of the alternatives suits their application needs best: permissionless voting does not rely on a central trusted service to prevent Sybil attacks, and is often implemented by providing voting rights proportional to wealth or computational power. However, it is known to centralise voting power in the hands of the wealthy. Permissioned voting introduces reliance on a centralised trusted authority to issue voting rights, or requires a practical and functional decentralised identity management solution. This permissioned alternative is often implemented through reputation based voting schemes, soulbound tokens, NFTs, or a Know Your Customer solution, checking the validity of participants’ identity.
6.3.2. 2. On-Chain Voting
One of many concerns of on-chain voting is the transaction fees associated with a vote. No DLT currently offers zero transaction fees, thus, when miners or users wish to vote, they must pay to do so. This further contributes to exacerbate inequalities as only the wealthier can afford to do so. These fees often soar unpredictably (Jain et al., 2023), and may also discriminate between voters as an election is occurring. The authors in (Park et al., 2021) have extensively documented many of the concerns of blockchain voting. Another concern pertaining to on-chain voting is the fact that censorship resistance cannot be assured (Wahrstätter et al., 2023): miners may simply refuse to approve blocks of certain users. Conversely, malicious agents can spam the network with empty blocks, thus delaying the approval of the user’s transactions. The feasibility of the latter highly depends on the access control mechanism of the DLT and the congestion control mechanism. Finally, current on-chain voting solutions do not offer ballot secrecy guarantees. The consequences arising from this are outlined in section 6.1.
It can therefore be observed that another compromise arises: binding elections are offered by on-chain voting, however, this means censorship resistance is not guaranteed, and participants must pay to vote.
6.4. Other vulnerabilities
There are other security concerns that arise when a combination of desirable properties are absent in the governance solution of the DLT.
6.4.1. 1. Bribe Automation
This exploit is a consequence of the following properties being absent: ballot secrecy, non-purchaseable voting power and fee-less voting. In DLTs where mining is competitive, if one mining pool gains a transaction, the other do not, so we can model the mining pools to be explicitly distinct from each other, and the smallest mining pool is of size 1 agent. Public votes make it evident to see how a mining pool voted by simply looking up the on-chain transactions. Less influential miners should either align their vote to those of the largest mining pools, or risk having their transactions end up in a fork and not receive the reward for the work they performed in including it. It is therefore easy for a briber to identify miners of less influence and bribe them to ‘vote’ for their desired protocol. The small sized mining pools will make a profit, and the briber can identify how many miners they must bribe to succeed in implementing their desired protocol. As the authors in (Austgen et al., 2023) showed, paradoxically, the more decentralised the network, the easier it is to create a systematic bribing cartel. Small sized miners are cheaper to bribe than large mining pools. Indeed, they even provide a protocol to automate this vote buying in DAOs. This can be re-purposed to buy a miner’s vote in a DLT. The bribe can be paid out as soon as the smart contract detects a transaction following the desired protocol of the briber.
6.4.2. 2. Submitting Malicious Protocol Updates:
Malicious protocol updates are not infrequent, and not only in DAOs or DLTs  [^9] Backdoor found in Library update, used by major Linux Distributions. Some go undetected due to low participation rates in governance, in others, the protocol update is forcibly approved by a whale, and in other occasions, the malicious protocol is hidden or not evidently detectable by the community. The consequences of this can be dire: double spending attacks, delaying governance outcomes, creating new tokens or loss of funds are amongst the many attack vectors possible.
It is not feasible to characterise all the ways in which malicious protocol updates may be submitted. This issue is application specific. Governance is only one of the components necessary in the DLT to prevent these exploits, but a robust governance protocol does not guarantee protection against malicious protocol updates. Access control mechanisms, security and trust assumptions are other components that must be taken into consideration. We encourage DLT design practitioners to make choices in their governance protocol that will alleviate the risk of this exploit. Namely, removing incentives to submit malicious proposals, or to comply with bribes to implement them.
7. Conclusion
We call for the DLT research and practitioners community to consider the matter of governance carefully. Current governance models make DLTs privy to a number of undesirable scenarios. Our strong encouragement is for practitioners to adopt well established notions of security in their own governance protocols, borrowing from the research in social choice theory, cryptography and e-voting schemes. We re-iterate notions of good governance properties, and present a collection of vulnerabilities that arise in the absence of them. We point practitioners towards exiting solutions that may serve as a starting point to begin to deliver good governance properties.
Acknowledgements.
Aida Manzano Kharman thanks IOTA Foundation for funding of her PhD studies. The authors thank Dr. Pietro Ferraro for his valuable advice.